# FPT approximations for Capacitated Sum of Radii and Diameters

### v1

@@ A.2 Notations and definitions
We assume that $k\leq\sqrt[4]{n/30}$ (and hence, $n\geq 30$), otherwise $n<30k^{4}$, and hence the brute force algorithm along with Lemma 3, running in FPT time in $k$ exactly solves uniform CapSoR. To see this, note that for CapSoR, we can guess the set of optimal centers and the corresponding radii correctly in time [^13] First guess the centers, there are (${n\choose k}$ options. Next for each center $x$, guess the furthest points $p$ assigned to $x$. There are $n$ options (and each such choice determines the corresponding cluster radius). Overall, there are (${n\choose k}n^{k}$ possible guesses. We can go over them one by one, checking for each guess whether it is feasible using Lemma 3, and thus find the optimum. (${n\choose k}n^{k+O(1)}$ which, in this case, is bounded by $2^{O(k\log k)}$.
@@ 20 Mark all $C_{i}\in\mathcal{C}_{F}$ such that $\Gamma^{-1}(C_{i})\neq\emptyset$;
22       ${\sf ball}(y_{i},r_{i})\leftarrow$ FIND-GOOD-DENSE-BALL$({\sf ball}(x_{i},2r_{i}),r_{i})$;
@@
Algorithm 11 FIND-GOOD-DENSE-BALL$({\sf ball}(x_{i},2r_{i}),r_{i})$
@@ A.4.1 Analysis
[…]
$>\sum_{t=1}^{4k}\left(\frac{n}{2k}-\frac{nt}{20k^{2}}\right)>\frac{3n}{10k}\cdot 4k>n.$
@@
$>\sum_{t=1}^{4k}\left(\frac{n}{2k}-\frac{nt}{20k^{2}}\right)>\frac{3n}{10k}\cdot 4k>n.$
@@
Intuition. Consider $a_{i}\in X\cup Z$ and let $C_{i}\in\mathcal{C}_{hit}$ hit by $a_{i}$. The key idea is that if $C_{i}$ is a regular cluster, then $a_{i}$ can serve all the points in $\mathcal{T}_{i}$ without violating the uniform capacity and within radius $2r_{i}+\ell(r_{i})$, since $|\mathcal{T}_{i}|\leq U$, as reasoned above. If $C_{i}$ is special, then it can happen that the number of points in $\mathcal{T}_{i}$ is larger than $U$, violating the uniform capacity. But, in this case, as mentioned above, there is an additional entry $(y_{i},r_{i})$ in $\mathcal{S}$. Further, 23 guarantees that ${\sf ball}(y_{i},r_{i})$ is a good dense ball for $C_{i}$, and hence we have $|{\sf ball}(y_{i},r_{i})\cap C_{i}|\geq\nicefrac{{n}}{{20k^{2}}}$. Therefore, in this case $y_{i}\in Y$ can serve $\nicefrac{{n}}{{20k^{2}}}$ points $D^{\prime}_{i}$ from $C_{i}$ within radius $r_{i}$, while $a_{i}$ can serve the remaining points in $\mathcal{T}_{i}$. Note that, $|\mathcal{T}_{i}\setminus D^{\prime}_{i}|\leq|\mathcal{T}_{i}|-|D^{\prime}_{i}|\leq|C_{i}|$. Since, $\mathcal{T}_{i}$’s form a partition of $P$, the total cost of this assignment is at most $3\sum_{i\in[k^{\prime}]}r_{i}\leq 3(1+\varepsilon){\rm OPT}$, as desired. Also, note that the number of centers opened is at most $k^{\prime}\leq k$.
@@
Note that $|\sigma^{\prime-1}(s_{i})|\leq|\mathcal{T}_{i}|\leq U$, since $C_{i}$ is a regular cluster. Also, ${\rm rad}(s_{i})\leq 2r_{i}+\ell(r_{i})$. Now, consider a special cluster $C_{i}$ and let $(s_{i},2r_{i}+\ell(r_{i}))$ and $(y_{i},r_{i})$ be the entries corresponding to $C_{i}$ in $\mathcal{S}$, where $s_{i}\in X$ and $y_{i}\in Y$. Recall that ${\sf ball}(y_{i},r_{i})$ returned by the subroutine FIND-GOOD-DENSE-BALL$({\sf ball}(s_{i},2r_{i}),r_{i})$ at Step 10 is a good dense ball for $C_{i}$, due to 23. Hence, we have $|{\sf ball}(y_{i},r_{i})\cap C_{i}|\geq\nicefrac{{n}}{{20k^{2}}}$. Let $D^{\prime}_{i}\subseteq D_{i}\cap C_{i}$ such that $|D^{\prime}_{i}|=\nicefrac{{n}}{{20k^{2}}}$. For $p\in\mathcal{T}_{i}$, define

### v2

@@ A.2 Notations and definitions
We assume that $k\leq\sqrt[4]{n/30}$ (and hence, $n\geq 30$), otherwise $n<30k^{4}$, and hence the brute force algorithm along with Lemma 3, running in FPT time in $k$ exactly solves uniform CapSoR. To see this, note that for CapSoR, we can guess the set of optimal centers and the corresponding radii correctly in time [^14] First guess the centers, there are $n\choose k$} options. Next for each center $x$, guess the furthest points $p$ assigned to $x$. There are $n$ options (and each such choice determines the corresponding cluster radius). Overall, there are ${n\choose k}n^{k}$ possible guesses. We can go over them one by one, checking for each guess whether it is feasible using Lemma 3, and thus find the optimum. ${n\choose k}n^{k+O(1)}$ which, in this case, is bounded by $2^{O(k\log k)}$.
@@ 20 Mark all $C_{i}\in\mathcal{C}_{F}$ such that $\Gamma^{-1}(C_{i})\neq\emptyset$;
[…]
@@
[…]
@@ A.4.1 Analysis
Let $S_{i_{j}}={\sf ball}(x_{i_{j}},2r_{i_{j}})$ and note that $C_{i_{j}}\subseteq S_{i_{j}}$. Also, let $F_{j}\subseteq F$ denote the points in $F$ when $C_{i_{j}}$ is processed at Step 10. Consider $j=1$ and note that $o_{i_{1}}\notin X\cup Z$ due to 3, and hence $o_{i_{1}}\in F_{1}$. Now, consider the execution of $\texttt{FIND-GOOD-DENSE-BALL}({\sf ball}(x_{i_{1}},r_{i_{1}}),r_{i_{1}})$ for the first $4k$ iterations. Let $d^{\ell}_{1}$ be the size of the densest ball within $S_{i_{1}}\cap P^{\prime\prime}$ centered at a point in $((S_{i_{1}}\cap F_{1})\setminus O)\cup\{o_{i_{1}}\}$ after iteration $\ell\leq 4k$. Then, note that $d^{\ell}_{1}>\max\{\frac{n}{2k}-\frac{n\ell}{20k^{2}},0\}\geq 1$, since the number of points remaining in $C_{i_{1}}$ after $\ell$ iterations is at least $\max\{\frac{n}{2k}-\frac{n\ell}{20k^{2}},0\}\geq 1$ and $o_{i}\in F_{1}$, due to 3. Thus, after every iteration $\ell\in[4k]$, the algorithm deletes at least ${\frac{n}{2k}-\frac{n\ell}{20k^{2}}}>\frac{3n}{10k}\geq 1$ many points from $S_{i_{1}}$. Hence, after $\ell=4k$ iterations, the total number of points deleted from $S_{i_{1}}$ is
$\sum_{t=1}^{4k}\left(\frac{n}{2k}-\frac{nt}{20k^{2}}\right)>\frac{3n}{10k}\cdot 4k>n.$
@@
$\sum_{t=1}^{4k}\left(\frac{n}{2k}-\frac{nt}{20k^{2}}\right)>\frac{3n}{10k}\cdot 4k>n.$
@@
Intuition. Consider $a_{i}\in X\cup Z$ and let $C_{i}\in\mathcal{C}_{hit}$ hit by $a_{i}$. The key idea is that if $C_{i}$ is a regular cluster, then $a_{i}$ can serve all the points in $\mathcal{T}_{i}$ without violating the uniform capacity and within radius $2r_{i}+\ell(r_{i})$, since $|\mathcal{T}_{i}|\leq U$, as reasoned above. If $C_{i}$ is special, then it can happen that the number of points in $\mathcal{T}_{i}$ is larger than $U$, violating the uniform capacity. But, in this case, as mentioned above, there is an additional entry $(y_{i},r_{i})$ in $\mathcal{S}$. Further, 16 guarantees that ${\sf ball}(y_{i},r_{i})$ is a good dense ball for $C_{i}$, and hence we have $|{\sf ball}(y_{i},r_{i})\cap C_{i}|\geq\nicefrac{{n}}{{20k^{2}}}$. Therefore, in this case $y_{i}\in Y$ can serve $\nicefrac{{n}}{{20k^{2}}}$ points $D^{\prime}_{i}$ from $C_{i}$ within radius $r_{i}$, while $a_{i}$ can serve the remaining points in $\mathcal{T}_{i}$. Note that, $|\mathcal{T}_{i}\setminus D^{\prime}_{i}|\leq|\mathcal{T}_{i}|-|D^{\prime}_{i}|\leq|C_{i}|$. Since, $\mathcal{T}_{i}$’s form a partition of $P$, the total cost of this assignment is at most $3\sum_{i\in[k^{\prime}]}r_{i}\leq 3(1+\varepsilon){\rm OPT}$, as desired. Also, note that the number of centers opened is at most $k^{\prime}\leq k$.
@@
[…]